# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 2.1 Conservation Equations for Reacting Air
We consider a model for reacting air comprising the five-species set $S=\{\mathrm{N_{2}},\mathrm{O_{2}},\mathrm{NO},\mathrm{N},\mathrm{O}\}$. The mass conservation equation for each species is:
@@
$w_{s}=\sum_{r\in R}w_{s,r}=M_{s}\sum_{r\in R}(\nu_{s,r}^{\prime\prime}-\nu_{s,r}^{\prime})\left(\mathcal{R}_{r}^{f}-\mathcal{R}_{r}^{b}\right),\qquad s\in S,$ (2)
where $M_{s}$ is the species molar mass and the sum extends over all reversible reactions $r$ in set $R$. Coefficients $\nu_{s,r}^{\prime}$ and $\nu_{s,r}^{\prime\prime}$ represent pre- and post-reaction stoichiometric coefficients for species $s$ in reaction $r$ written in generic form, when read left-to-right:
$\sum_{q\in S}\nu_{q,r}^{\prime}\,\mathcal{A}_{q}\xrightleftharpoons[k_{r}^{b}]{k_{r}^{f}}\sum_{s\in S}\nu_{s,r}^{\prime\prime}\,\mathcal{A}_{s},\qquad r\in R.$ (3)
Here, $\mathcal{A}_{q}$ and $\mathcal{A}_{s}$ represent the chemical symbol of a particular species on the reactant, or product side of the equation. If a given species does not participate as reactant or product, its corresponding pre- or post-reaction stoichiometric coefficient in Eq. (3) is zero. The terms $\mathcal{R}_{r}^{f}$ and $\mathcal{R}_{r}^{b}$ in Eq. (2) represent the forward and backward rates of reaction $r$ per unit volume respectively:
$\mathcal{R}_{r}^{f}=k_{r}^{f}\prod_{q\in S}\left(\frac{\rho_{q}}{M_{q}}\right)^{\nu_{q,r}^{\prime}},\qquad\mathcal{R}_{r}^{b}=k_{r}^{b}\prod_{s\in S}\left(\frac{\rho_{s}}{M_{s}}\right)^{\nu_{s,r}^{\prime\prime}},\qquad r\in R.$ (4)
@@
In this work we assume that the vibrational energy of the gas mixture can be represented with a single vibrational energy conservation equation, characterized by a mixture vibrational temperature $T_{v}$:
$\frac{\partial E_{v}}{\partial t}+\nabla\cdot(E_{v}\vec{u})+\nabla\cdot\sum_{s\in S}(\rho_{s}\,e_{v,s}\vec{v}_{s})+\nabla\cdot\vec{q}_{v}=Q_{v}^{\mathrm{relax}}+Q_{v}^{\mathrm{chem}}.$ (5)
@@
Terms on the left hand side account for the local rate of change in $E_{v}$ and vibrational energy transport, while the two vibrational energy source terms sit on the right hand side. The first, $Q_{v}^{\mathrm{relax}}$, represents the vibration-translation relaxation source term. It is typically modeled as a sum of Landau-Teller terms:
$Q_{v}^{\mathrm{relax}}=\sum_{s\in D}\rho_{s}\frac{e_{v,s}(T)-e_{v,s}(T_{v})}{\langle\tau_{s}^{\mathrm{v}}\rangle},$ (6)
@@
where $D=\{\mathrm{N_{2}},\mathrm{O_{2}},\mathrm{NO}\}$ represents the set of diatomic species contributing to the vibrational energy mode and $\langle\tau_{s}^{\mathrm{v}}\rangle$ is an appropriately averaged vibrational relaxation time for species $s$. Refer to Sec. 2.3 for further details on how we compute vibrational relaxation times within the context of the current multi-temperature model.
@@
The chemistry-vibrational energy coupling term $Q_{v}^{\mathrm{chem}}=\sum\limits_{s\in D}\sum\limits_{r\in R}Q_{v,r,s}^{\mathrm{chem}}$ in Eq. (5) accounts for net removal, or replenishment of vibrational energy during chemical reactions. When written in generic form, the individual terms may be expressed as:
@@
$Q_{v,r,s}^{\mathrm{chem}}=(\nu_{s,r}^{\prime\prime}-\nu_{s,r}^{\prime})\left(\langle\varepsilon_{v,s}\rangle_{r}^{f}\,\mathcal{R}_{r}^{f}-\langle\varepsilon_{v,s}\rangle_{r}^{b}\,\mathcal{R}_{r}^{b}\right)\qquad r\in R,s\in D$ (7)
@@
where $\langle\varepsilon_{v,s}\rangle_{r}^{f}$ represents the average amount of vibrational energy per unit mole of species $s$ being removed in the forward sense of reaction $r$, whereas $\langle\varepsilon_{v,s}\rangle_{r}^{b}$ is the corresponding amount being replenished by the reverse process. In principle, $\langle\varepsilon_{v,s}\rangle_{r}^{f}$ and $\langle\varepsilon_{v,s}\rangle_{r}^{b}$ may take on different values, for instance when the local gas state exhibits significant departure from thermal equilibrium. However, when $\langle\varepsilon_{v,s}\rangle_{r}^{f}\approx\langle\varepsilon_{v,s}\rangle_{r}^{b}=\langle\varepsilon_{v,s}\rangle_{r}$ holds, Eq. (7) can be simplified to:
@@
$Q_{v,r,s}^{\mathrm{chem}}=(\nu_{s,r}^{\prime\prime}-\nu_{s,r}^{\prime})\left(\mathcal{R}_{r}^{f}-\mathcal{R}_{r}^{b}\right)\langle\varepsilon_{v,s}\rangle_{r}\qquad r\in R,s\in D$ (8)
@@
Closure of Eq. (5) requires all relevant $\langle\varepsilon_{v,s}\rangle_{r}$ also be provided by the chemical kinetics model. One common option is to assume non-preferential vibrational energy-chemistry coupling. This assumes that the amount of vibrational energy of species $s$ removed/replenished per reaction $r$ will be equal to the average of that species in the mixture at the local thermodynamic state, i.e. $\langle\varepsilon_{v,s}\rangle_{r}=e_{v,s}(T_{v})/M_{s}$. When substituted into Eq. (8) for every species and combined with Eq. (2), this yields the non-preferential (i.e. vibrationally unbiased) form of the overall source term:
$Q_{v}^{\mathrm{chem,non-pref}}=\sum_{s\in D}w_{s}\,e_{v,s}(T_{v}).$ (9)
@@
It is revealed by the QCT calculations in Secs. 3 and 4 that assuming non-preferential vibrational energy removal is a poor fit for dissociation reactions. Such reactions exhibit significant vibrational bias in their reaction rate coefficients, as well as for the average vibrational energy removed by dissociation. Thus, a major feature of the chemical-kinetics model described in Sec. 2.2 is that it accounts for this preferential treatment by providing a more accurate expression for $\langle\varepsilon_{v,s}\rangle_{r}$ in dissociation reactions.
2.2 The MMT Model
@@
In this work we propose modifications to the classic Marrone-Treanor model to represent air dissociation dynamics. The Modified Marrone-Treanor model is motivated by the analysis of a large quasi-classical trajectory (QCT) database created using accurate potential energy surfaces developed by researchers at the University of Minnesota [31]. At its core the MMT model consists of three main parts: (1) an analytical expression for the two-temperature nonequilibrium dissociation rate coefficient, (2) a corresponding expression for the vibrational energy change per dissociation and (3) a set of QCT- and DMS-derived kinetic parameters necessary to compute both these quantities. The MMT model’s nonequilibrium dissociation rate coefficient $k_{d}^{\mathrm{MMT-NB}}$ is calculated as:
$k_{d}^{\mathrm{MMT-NB}}(T,T_{v})=k_{d}^{\mathrm{arr}}\,Z(T,T_{v})\,f_{k}^{\mathrm{NB}},$ (10)
@@
where $k_{d}^{\mathrm{arr}}$ represents the thermal equilibrium dissociation rate coefficient in modified Arrhenius form:
$k_{d}^{\mathrm{arr}}(T)=CT^{n}\exp(-T_{D,s}/T),$ (11)
@@
while $Z(T,T_{v})$ is the MMT model’s vibrational nonequilibrium factor, exactly as derived by Marrone and Treanor [5]:
$Z(T,T_{v})=\frac{Q(T)\,Q(T_{F})}{Q(T_{v})\,Q(-U)},$ (12)
@@
and $f_{k}^{\mathrm{NB}}$ is a non-Boltzmann correction discussed in more detail in Secs. 5 and 6. The $Q^{\prime}s$ appearing in Eq. (12) are approximate vibrational partition functions of the form:
$Q(T)=\frac{1-\exp(-T_{D,s}/T)}{1-\exp(-\theta_{v,s}/T)}$ (13)
and $T_{F}$ and $U$ are pseudo-temperatures defined as:
$T_{F}(T,T_{v})=\left(\frac{1}{T_{v}}-\frac{1}{T}-\frac{1}{U(T)}\right)^{-1}$ (14)
and:
@@
respectively. Thus, in addition to the local $(T,T_{v})$-values, Eqs. (10)-(15) depend on the reaction-specific Arrhenius parameters $C$ and $n$, characteristic vibrational and dissociation temperatures $\theta_{v,s}$ and $T_{D,s}$ of the dissociating diatom, as well as the MMT reaction-specific parameters $U^{*}$ and $a_{U}$.
@@
$\langle\varepsilon_{v,s}\rangle_{d}^{\mathrm{Knab}}(T,T_{v})=\frac{\mathrm{k_{B}}\,\theta_{v,s}}{\exp(\theta_{v,s}/T_{F})-1}-\frac{\mathrm{k_{B}}\,T_{D,s}}{\exp(T_{D,s}/T_{F})-1},$ (16)
where, by convention, $\langle\varepsilon_{v,s}\rangle_{d}^{\mathrm{Knab}}>0$ when vibrational energy is being removed during dissociation of species $s$. Apart from being parameterized by the dissociating species’ characteristic vibrational and dissociation temperatures, this expression shares with the MMT rate coefficient an indirect dependence on vibrational temperature through the pseudo-temperature $T_{F}(T,T_{v})$. As discussed in Sec. 4.4, we take advantage of this commonality to determine the MMT parameters $a_{U}$ and $U^{*}$ applicable to both the vibrational energy removed and the rate coefficient from QCT data for $\langle\varepsilon_{v,s}\rangle_{d}$ under thermal-equilibrium conditions.
@@
$\langle\varepsilon_{v,s}\rangle_{d}^{\mathrm{MMT-NB}}=\langle\varepsilon_{v,s}\rangle_{d}^{\mathrm{Knab}}f_{\varepsilon}^{\mathrm{NB}}.$ (17)
Table 2 lists model parameters for $\mathrm{N_{2}}$ dissociation with collision partners $\mathrm{M}=\{\mathrm{N_{2}},\mathrm{N},\mathrm{O_{2}}\}$ and for $\mathrm{O_{2}}$ dissociation with partners $\mathrm{M}=\{\mathrm{O_{2}},\mathrm{O},\mathrm{N_{2}}\}$ respectively, originally reported in Tables 3.3 and 3.4 of Ref. [18]. When curve-fitting the QCT rate coefficients to Eq. (11) the dissociation temperature was treated as a third unknown in addition to $C$ and $n$. Thus, $T_{D,\mathrm{N_{2}}}$ and $T_{D,\mathrm{O_{2}}}$ differ for every reaction listed in Table 2, even for reactions where the dissociating species $s$ is the same. The MMT model parameters listed in Table 2 are used in this article to demonstrate the accuracy of the model formulation and describe its implementation for several reactions of interest. It is important to note that a subsequent companion article will present a more complete set of MMT parameters for 5-species air (with some differences) based on a more recent and extensive set of PESs.

### v2

@@ 2.1 Conservation equations for a reacting gas mixture
We consider a model for a reacting gas mixture in thermal and chemical nonequilibrium, where the generic species set $S$ comprises diatomic and atomic species (for 5-species air $S=\{\mathrm{N_{2}},\mathrm{O_{2}},\mathrm{NO},\mathrm{N},\mathrm{O}\}$, but for the calculations in Sec. 7, either $S=\{\mathrm{N_{2}},\mathrm{N}\}$, or $S=\{\mathrm{O_{2}},\mathrm{O}\}$). The mass conservation equation for each species is:
@@
$w_{s}=\sum_{r\in R}w_{s,r}=M_{s}\sum_{r\in R}(\nu_{s,r}^{b}-\nu_{s,r}^{f})\left(\mathcal{R}_{r}^{f}-\mathcal{R}_{r}^{b}\right),\qquad s\in S,$ (2)
where $M_{s}$ is the species molar mass and the sum extends over all reversible reactions $r$ in set $R$. Coefficients $\nu_{s,r}^{f}$ and $\nu_{s,r}^{b}$ represent pre- and post-reaction stoichiometric coefficients for species $s$ in reaction $r$ written in generic form, when read left-to-right:
$\sum_{q\in S}\nu_{q,r}^{f}\,\mathcal{A}_{q}\xrightleftharpoons[k_{r}^{b}]{k_{r}^{f}}\sum_{s\in S}\nu_{s,r}^{b}\,\mathcal{A}_{s},\qquad r\in R.$ (3)
Here $\mathcal{A}_{q}$ and $\mathcal{A}_{s}$ represent the chemical symbol of a particular species on the reactant, or product side of the equation. If a given species does not participate as reactant or product, its corresponding pre- or post-reaction stoichiometric coefficient in Eq. (3) is zero. The terms $\mathcal{R}_{r}^{f}$ and $\mathcal{R}_{r}^{b}$ in Eq. (2) represent the forward and backward rates of reaction $r$ per unit volume respectively:
$\mathcal{R}_{r}^{f}=k_{r}^{f}\prod_{q\in S}\left(\frac{\rho_{q}}{M_{q}}\right)^{\nu_{q,r}^{f}},\qquad\mathcal{R}_{r}^{b}=k_{r}^{b}\prod_{s\in S}\left(\frac{\rho_{s}}{M_{s}}\right)^{\nu_{s,r}^{b}},\qquad r\in R.$ (4)
@@
In this work we assume for simplicity that the vibrational energy density of the gas mixture $E_{\mathrm{v}}$ can be represented with a single conservation equation, characterized by a mixture vibrational temperature $T_{\mathrm{v}}$:
$\frac{\partial E_{\mathrm{v}}}{\partial t}+\nabla\cdot(E_{\mathrm{v}}\vec{u})+\nabla\cdot\sum_{s\in D}(\rho_{s}\,e_{\mathrm{v},s}\vec{v}_{s})+\nabla\cdot\vec{q}_{\mathrm{v}}=w_{\mathrm{v}}^{\mathrm{relax}}+w_{\mathrm{v}}^{\mathrm{chem}},$ (5)
where $D$ represents the set of diatomic species contributing to the vibrational energy mode. In a mixture containing multiple diatomic species, such as five-species air, this would be $D=\{\mathrm{N_{2}},\mathrm{O_{2}},\mathrm{NO}\}$ and $T_{\mathrm{v}}$ would be implicitly defined through the relation $E_{\mathrm{v}}=\sum\limits_{s\in D}\rho_{s}\,e_{\mathrm{v},s}(T_{\mathrm{v}})$. In the simpler diatom/atom ($\mathrm{A_{2}/A}$) mixtures considered in Sec. 7, this automatically reduces to $D=\{\mathrm{A_{2}}\}$ and implies that $T_{\mathrm{v}}=T_{\mathrm{v,A_{2}}}$.
@@
Terms on the left hand side of Eq. (5) account for the local rate of change in $E_{\mathrm{v}}$ and vibrational energy transport by advection and diffusion respectively, while the two vibrational energy source terms sit on the right hand side. The first, $w_{\mathrm{v}}^{\mathrm{relax}}$, represents the vibration-translation relaxation source term. It is typically modeled as a sum of Landau-Teller terms:
$w_{\mathrm{v}}^{\mathrm{relax}}=\sum_{s\in D}\rho_{s}\frac{e_{\mathrm{v},s}(T)-e_{\mathrm{v},s}(T_{\mathrm{v}})}{\langle\tau_{s}^{\mathrm{v}}\rangle},$ (6)
@@
where $\langle\tau_{s}^{\mathrm{v}}\rangle$ is an appropriately averaged vibrational relaxation time for species $s$. Refer to Sec. 2.3 for further details on how we compute vibrational relaxation times within the context of the current multi-temperature model.
@@
The chemistry-vibrational energy coupling term $w_{\mathrm{v}}^{\mathrm{chem}}=\sum\limits_{s\in D}\sum\limits_{r\in R}w_{\mathrm{v},r,s}^{\mathrm{chem}}$ in Eq. (5) accounts for net removal, or replenishment of vibrational energy during chemical reactions. When written in generic form, the individual terms may be expressed as:
@@
$w_{\mathrm{v},r,s}^{\mathrm{chem}}=(\nu_{s,r}^{b}-\nu_{s,r}^{f})\left(\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{f}\,\mathcal{R}_{r}^{f}-\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{b}\,\mathcal{R}_{r}^{b}\right)\qquad r\in R,s\in D$ (7)
@@
where $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{f}$ represents the average amount of vibrational energy per unit mole of species $s$ being removed in the forward sense of reaction $r$, whereas $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{b}$ is the corresponding amount being replenished by the reverse process. In principle, $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{f}$ and $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{b}$ may take on different values, for instance when the local gas state exhibits significant departure from thermal equilibrium. However, when $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{f}\approx\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{b}=\langle\varepsilon_{\mathrm{v},s}\rangle_{r}$ holds, Eq. (7) can be simplified to:
@@
$w_{\mathrm{v},r,s}^{\mathrm{chem}}=(\nu_{s,r}^{b}-\nu_{s,r}^{f})\left(\mathcal{R}_{r}^{f}-\mathcal{R}_{r}^{b}\right)\langle\varepsilon_{\mathrm{v},s}\rangle_{r}\qquad r\in R,s\in D$ (8)
@@
Closure of Eq. (5) requires all relevant $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}$ also be provided by the chemical kinetics model. One common option is to assume non-preferential vibrational energy-chemistry coupling. This assumes that the amount of vibrational energy of species $s$ removed/replenished per reaction $r$ will be equal to the average of that species in the mixture at the local thermodynamic state, i.e. $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}^{\mathrm{non-pref.}}=e_{\mathrm{v},s}(T_{\mathrm{v}})/M_{s}$. When substituted into Eq. (8) for every species and combined with Eq. (2), this yields the non-preferential (i.e. vibrationally unbiased) form of the overall source term:
$w_{\mathrm{v}}^{\mathrm{chem,(non-pref.)}}=\sum_{s\in D}w_{s}\,e_{\mathrm{v},s}(T_{\mathrm{v}}).$ (9)
@@
It is revealed by the QCT calculations in Secs. 3 and 4 that assuming non-preferential vibrational energy removal is a poor fit for dissociation reactions. Such reactions exhibit significant vibrational bias in their reaction rate coefficients, as well as for the average vibrational energy removed by dissociation. Thus, a major feature of the chemical-kinetics model described in Sec. 2.2 is that it accounts for this preferential treatment by providing a more accurate expression for $\langle\varepsilon_{\mathrm{v},s}\rangle_{r}$ in dissociation reactions.
2.2 The MMT model formulation
@@
In this work we propose modifications to the classic Marrone-Treanor model to represent air dissociation dynamics. The Modified Marrone-Treanor model is motivated by the analysis of a large quasiclassical trajectory database created using accurate potential energy surfaces developed by researchers at the University of Minnesota [38]. At its core the MMT model consists of three main parts: (1) an analytical expression for the two-temperature nonequilibrium dissociation rate coefficient, (2) a corresponding expression for the vibrational energy change per dissociation and (3) a set of QCT- and DMS-derived kinetic parameters necessary to compute both these quantities. The MMT model’s non-Boltzmann, two-temperature dissociation rate coefficient $k_{\mathrm{diss}}^{\mathrm{MMT-NB}}$ is defined as:
$k_{\mathrm{diss}}^{\mathrm{MMT-NB}}(T,T_{\mathrm{v}})=k^{\mathrm{Arr}}(T)\,Z(T,T_{\mathrm{v}})\,f_{k}^{\mathrm{NB}},$ (10)
@@
where $k^{\mathrm{Arr}}$ represents the thermal equilibrium dissociation rate coefficient in modified Arrhenius form:
$k^{\mathrm{Arr}}(T)=CT^{n}\exp(-T_{\mathrm{D},s}/T),$ (11)
@@
while $Z(T,T_{\mathrm{v}})$ is the MMT model’s vibrational nonequilibrium factor, exactly as derived by Marrone and Treanor [5]:
$Z(T,T_{\mathrm{v}})=\frac{\tilde{Q}_{\mathrm{v},s}(T)\,\tilde{Q}_{\mathrm{v},s}(T_{F})}{\tilde{Q}_{\mathrm{v},s}(T_{\mathrm{v}})\,\tilde{Q}_{\mathrm{v},s}(-U)},$ (12)
@@
and $f_{k}^{\mathrm{NB}}$ is a non-Boltzmann correction discussed in more detail in Secs. 5 and 6. The $\tilde{Q}$’s appearing in Eq. (12) are approximate vibrational partition functions of the form:
$\tilde{Q}_{\mathrm{v},s}(T)=\frac{1-\exp\,(-T_{\mathrm{D},s}/T)}{1-\exp\,(-\Theta_{\mathrm{v},s}/T)}$ (13)
and $T_{F}$ and $U$ are pseudotemperatures defined as
$T_{F}(T,T_{\mathrm{v}})=\left(\frac{1}{T_{\mathrm{v}}}-\frac{1}{T}-\frac{1}{U(T)}\right)^{-1}$ (14)
and
@@
respectively. Thus, in addition to the local $(T,T_{\mathrm{v}})$ values, Eqs. (10)-(15) depend on the reaction-specific Arrhenius parameters $C$ and $n$, characteristic vibrational and dissociation temperatures $\Theta_{\mathrm{v},s}$ and $T_{\mathrm{D},s}$ of the dissociating diatom, as well as the MMT reaction-specific parameters $U^{*}$ and $a_{U}$.
@@
$\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}^{\mathrm{Knab}}(T,T_{\mathrm{v}})=\frac{\mathrm{k_{B}}\,\Theta_{\mathrm{v},s}}{\exp\,(\Theta_{\mathrm{v},s}/T_{F})-1}-\frac{\mathrm{k_{B}}\,T_{\mathrm{D},s}}{\exp\,(T_{\mathrm{D},s}/T_{F})-1},$ (16)
where, by convention, $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}^{\mathrm{Knab}}>0$ when vibrational energy is being removed during dissociation of species $s$. Apart from being parameterized by the dissociating species’ characteristic vibrational and dissociation temperatures, this expression shares with the MMT rate coefficient an indirect dependence on vibrational temperature through the pseudotemperature $T_{F}(T,T_{\mathrm{v}})$. As discussed in Sec. 4.4, we take advantage of this commonality to determine the MMT parameters $a_{U}$ and $U^{*}$ applicable to both the vibrational energy removed and the rate coefficient from QCT data for $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$ under thermal-equilibrium conditions, i.e. where $T=T_{\mathrm{v}}$ and $f_{k}^{\mathrm{NB}}=f_{\varepsilon}^{\mathrm{NB}}=1$.
@@
$\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}^{\mathrm{MMT-NB}}=\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}^{\mathrm{Knab}}f_{\varepsilon}^{\mathrm{NB}}.$ (17)
In Table 2 we list model parameters for $\mathrm{N_{2}}$ dissociation with collision partners $\mathrm{M}=\{\mathrm{N_{2}},\mathrm{N},\mathrm{O_{2}}\}$ and for $\mathrm{O_{2}}$ dissociation with partners $\mathrm{M}=\{\mathrm{O_{2}},\mathrm{O},\mathrm{N_{2}}\}$ respectively. We present two distinct, but nearly equivalent parameter sets. In sub-table 2(a) we list values obtained when the QCT-derived rate coefficients are curve-fit to Eq. (11) with the dissociation temperature treated as the third unknown in addition to $C$ and $n$. Thus, $T_{\mathrm{D},\mathrm{N_{2}}}$ and $T_{\mathrm{D},\mathrm{O_{2}}}$ differ for every reaction listed in Table 2(a), even for reactions where the dissociating diatomic species $\mathrm{A_{2}}$ is the same. In sub-table 2(b) we list values obtained from alternative two-parameter curve fits with dissociation temperatures held fixed at the commonly accepted values $T_{\mathrm{D},\mathrm{N_{2}}}=113\,200\mathrm{K}$ and $T_{\mathrm{D},\mathrm{O_{2}}}=59\,380\mathrm{K}$ respectively. The values for $C$, $n$ and $T_{\mathrm{D},\mathrm{A_{2}}}$ (columns 3-5) appeared originally in Tables 3.1 and 3.2 of Ref. [25] on rows marked as “effective dissociation” (i.e. sum of simple + swap dissociation [^1] Simple, or single dissociation refers to quasiclassical trajectories of type $\mathrm{AB}+\mathrm{CD}\rightarrow\mathrm{A}+\mathrm{B}+\mathrm{CD}$, as opposed to those involving simultaneous swap-dissociation $\mathrm{AB}+\mathrm{CD}\rightarrow\mathrm{A}+\mathrm{C}+\mathrm{BD}$, or double-dissociation $\mathrm{AB}+\mathrm{CD}\rightarrow\mathrm{A}+\mathrm{B}+\mathrm{C}+\mathrm{D}$. rates). Differences in MMT parameters $a_{U}$ and $U^{*}$ (columns 7-8) between sub-tables 2(a) and 2(b) for the same reaction are a consequence of curve-fitting the QCT data for vibrational energy change per dissociation to Eq. (16) with different values of $T_{\mathrm{D},\mathrm{A_{2}}}$ (see Sec. 4.4). These parameters were originally listed in Tables 3.3 and 3.4 of Ref. [25].
Further note minor differences in the listed values for characteristic vibrational temperature of the dissociating diatomic species (column 6) both in Tables 2(a) and 2(b). Whereas $\Theta_{\mathrm{v},\mathrm{N_{2}}}=3\,411\,\mathrm{K}$ for $\mathrm{N_{2}}$ dissociation with partners $\mathrm{N_{2}}$ and $\mathrm{N}$ (in rows 1-2), it becomes $\Theta_{\mathrm{v},\mathrm{N_{2}}}=3\,415\,\mathrm{K}$ for dissociation with collision partner $\mathrm{O_{2}}$ (3rd row). The analog happens for $\mathrm{O_{2}}$ dissociation with collision partners $\mathrm{O_{2}}$ and $\mathrm{O}$ (rows 4-5), where $\Theta_{\mathrm{v},\mathrm{O_{2}}}=2\,280\,\mathrm{K}$ vs. $2\,263\,\mathrm{K}$ with partner $\mathrm{N_{2}}$ (row 6). Such discrepancies can be traced back to the slightly different shapes of diatomic potential curves $V_{D}(r_{\mathrm{A-A}})$ associated with different PES versions. The characteristic vibrational temperatures listed in Table 2 are defined as $\Theta_{\mathrm{v},\mathrm{A_{2}}}=[\varepsilon_{\mathrm{A_{2}}}(v\!=\!1)-\varepsilon_{\mathrm{A_{2}}}(v\!=\!0)]/\mathrm{k_{B}}$ and thus depend on the precise vibrational energies of the lowest-energy level and its immediate higher-lying neighbor. The values of all $\varepsilon_{\mathrm{A_{2}}}(v)$, in turn, are found as approximate solutions to an eigenvalue problem taking $V_{D}(r_{\mathrm{A-A}})$ as an input (see Sec. 3.2 of Bender [39] for more details). Whereas all trajectory calculations for reactions in rows 1-2 were performed on the same $\mathrm{N_{4}}(^{1}A)$ PES [13, 14, 39], QCT calculations for reaction 3 relied on the $\mathrm{N_{2}O_{2}}(^{3}A)$ PES [20] instead. The latter PES employed slightly different shape parameters in defining the nitrogen diatomic potential $V_{D}(r_{\mathrm{N-N}})$. In a similar fashion, reactions 4-5 employed the combined set of $\mathrm{O_{4}}$ [17, 18] and $\mathrm{O_{3}}$ [19] PESs all sharing a common diatomic potential curve for $\mathrm{O_{2}}$. Reaction in row 6 again relied on the triplet $\mathrm{N_{2}O_{2}}$ PES, whose shape parameters for defining $V_{D}(r_{\mathrm{O-O}})$ differed slightly from those for the pure oxygen PESs.
At the fluid scale of the two-temperature model proposed in this work, the slight discrepancies in vibrational energy ladders generated with different PESs have a negligible influence on the overall predictions and can be mostly ignored. By contrast, in multi-species DMS calculations [34, 37] that resolve individual molecular collisions, a common set of diatomic potential curves across all PESs involved becomes crucial to maintaining energy conservation over successive time steps. This consideration prompted the Minnesota chemists to recently re-fit all their currently available PESs with a consistent set of $V_{D}(r_{\mathrm{N-N}})$, $V_{D}(r_{\mathrm{O-O}})$ and $V_{D}(r_{\mathrm{N-O}})$ curves. In an upcoming article we rely on this updated set of PESs to generate new MMT model parameters applicable to a five-species air mixture ($S=\{\mathrm{N_{2}},\mathrm{O_{2}},\mathrm{NO},\mathrm{N},\mathrm{O}\}$). Regarding the present work, the parameters of Table 2 are sufficient for benchmarking the MMT model against DMS reference calculations of simpler $\mathrm{N_{2}/N}$ and $\mathrm{O_{2}/O}$ mixtures unaffected by the aforementioned inconsistencies (see Sec. 7).
Columns 9-10 of Table 2 list the non-Boltzmann correction factors applied in Eqs. (10) and (17) respectively. Refer to Sec. 5 for further details and the justification for employing common values of $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ for all six reactions. In summary, Table 2 contains all parameters to completely evaluate the MMT dissociation rate coefficient and vibrational energy loss terms, regardless of whether one prefers those derived from variable, or fixed-$T_{\mathrm{D}}$ Arrhenius fits.
Finally, recall that Eqs. (10)-(17) have been implicitly formulated within the framework of a fluid model with only a single vibrational temperature. Specifically, wherever a vibrational temperature is required as input to the MMT expressions, the mixture-averaged value $T_{\mathrm{v}}$ is used. However, if one was to implement the model within a multi-$T_{\mathrm{v}}$ fluid description, one should replace $T_{\mathrm{v}}$ everywhere in Eqs. (10)-(17) with the vibrational temperature of the particular dissociating species.